OpenStudy (anonymous):

I roll a cube two times. Find p (even, then not 2) I have to write the probability as a fraction @Zale101

OpenStudy (ja1):

@zarkon not zurkon :)

OpenStudy (anonymous):

oh ahha thanks @Zarkon

OpenStudy (anonymous):

Can you help @JA1

OpenStudy (ja1):

Sorry probability and me don't match :3

OpenStudy (anonymous):

ok thanks I am trying to understand @JA1

OpenStudy (ja1):

No prob :)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

\[P_{even} = \frac{1}{2}\]\[P_{not2} = \frac{5}{6}\]\[P_{even,thennot2} = P_{even}*P_{not2} = \frac{5}{12}\] Alternative route: count number of green leaves in tree diagram, divide by total number of green + red leaves.